OpenStudy (anonymous):

Compound functions: f(x)=2x^2-5 and g(x)=x^2-3x+1 Find: (f+g)(1) I got 3x^2-3x-4. Is that right?? Find: (f-g)(-2) How do I even do this one? I got-2x^2-6xx+12 Find: (f*g)(4) I have no idea but I got 8x^4-24x^3-12x^2+60x-20 Find: (f/g)(1.5) I don't even know where to begin with this one. Do I use polynomial division or something?

OpenStudy (anonymous):

That's either Advanced 8th grade math, or Freshman year Algebra...

OpenStudy (anonymous):

\[(f+g)(x)=f(x)+g(x)\]so\[(f+g)(1)= f(1)+g(1)\]For f(1) plug in 1 in the place of x in the expression given for f(x)\[f(1)= 2\times1^2-5=-3\]Similarly solve for g(1) and add for (f+g)(1). For the other parts follow similar steps using the following relations: \[(f-g)(x)= f(x)-g(x)\]\[(f*g)(x)=f(x)*g(x)\]\[(f/g)(x) = \frac{f(x)}{g(x)}\]

OpenStudy (anonymous):

WAIT WHAT!??? Sorry it took a while for me to reply. You gotta be kidding me. I did all of these questions wrong lol.

OpenStudy (anonymous):

This is also precalculus lol chasingthepursuit :P